MIDC/Data Quality Statements
Total Records: 19, displaying 1-19

R#SiteSerial NumberMeasurementStart DateStart TimeEnd DateEnd TimeData Quality Statement
1BMSDirect LI-20112/21/201307:00:0012/21/201316:22:00Due to a scheduled NREL power outage, to upgrade campus electrical infrastructure, The following BMS instruments were not pointing at the sun during this time: Direct NIP #1, LI-201, LI-190, NIP RG780, TWC; and the following instruments were un-shaded: Diffuse CM22, CUV4, CG4. The Middle Sci-Tec tracker for these instruments were temporarily put on a secondary UPS which failed immediately.
2BMSDirect LI-20112/16/201309:00:00The Direct LI-201 sensor has been temporarily moved (swapped with the LI-190) to to the outer "piggy back" lower position on the tracker, so that the Direct Quantum LI-190 sensor would not be shaded by the ventilator on the tracker to the south of the instrument. Until it gets moved to the final upper position, the sensor will be shaded between approx 9:10 and 9:40 MST of every day.
3BMSDirect LI-20103/02/201313:06:0003/02/201318:24:00Due to a scheduled NREL power outage, to upgrade campus power capacity, The following BMS instruments were not pointing at the sun during this time: Direct CH1, NIP #2, NIP #1, LI-201, NIP RG780, CUVA2, CUVB2, TWC; and the following instruments were un-shaded: Diffuse 8-48, PSP, CM22, CUV4, PIR, CG4. The Sci-Tec tracker for these insruments are protected by UPS, however the battery depleted at 13:06 MST (the power outage started at 8:06 MST)
4BMSDirect LI-20105/26/201206:15:0005/26/201207:12:00Due to a planned power outage for SRRL electrical work, the following BMS instruments were not following the sun: NIP #1, RG780 NIP, LI-201, TWC Photometer and the following instruments were not shaded CM22, CUV4, and CG4.
5BMSDirect LI-20109/08/201108:23:0009/08/201109:38:00Due to a planned power outage for SRRL construction activities, the following BMS instruments were not following the sun: NIP #1, RG780 NIP, LI-201, TWC Photometer and the following instruments were not shaded CM22, CUV4, and CG4.
6BMSDirect LI-20105/07/201105:55:0005/09/201107:58:00There was a planned power outage for SRRL construction activities that started at 5:18 MST. Unfortunately, the backup generator provided to SRRL was providing bad power to the UPS powering the Sci-Tec trackers, therefore the UPS failed at 5:55 MST. Unfortunately, when normal AC power was restored at 11:45 MST, the Sci-Tec trackers had lost all their configuration due to UPS cycling on and off during the outage, therefore were not tracking. The tracker settings were restored on May 9 at 7:58 MST. The following BMS instruments were not pointing at the sun during this time: Direct CH1, NIP #2, NIP #1, LI-201, NIP RG780, CUVA2, CUVB2, TWC; and the following instruments were un-shaded: Diffuse 8-48, PSP, CM22, CUV4, PIR, CG4
7BMSDirect LI-20104/15/201106:43:0004/15/201106:50:00There was a planned power outage for SRRL construction activities that started at 6:11 MST. Unfortunately, the UPS powering the Sci-Tec trackers failed due to bad/old batteries at 6:43 MST. The following BMS instruments were not point at the sun for this brief period of time: CH1, NIP #2, NIP #1, LI-201, NIP RG780.
8BMSDirect LI-20103/18/201100:00:0003/18/201107:57:00The Sci-Tec #2 tracker was jammed on the shading plate. The following instruments were unshaded: CM22, CUV4, CG4; and the following instruments were not pointing at the sun: NIP #1, NIP RG780, LI-201, TWC Photometer.
9BMSPYHR100Direct LI-20110/30/201000:00:0011/01/201011:27:00The Sci-Tec #2 tracker was apparently not working during this time. The following instruments were unshaded: CM22, CUV4, CG4; and the following instruments were not pointing at the sun: NIP #1, NIP RG780, LI-201.
10BMSDirect LI-20112/15/200909:40:0012/15/200910:00:00The Sci-Tec #2 tracker collided with another empty Sci-Tec tracker. The following instruments were unshaded: CM22, CUV4, CG4; and the following instruments were not pointing at the sun: NIP #1, NIP RG780, LI-201.
11BMSDirect LI-20112/09/200912:11:0012/09/200914:22:00Due to a power outage at NREL, followed by bad power, the SRRL tracker UPS shutdown. The following BMS instruments were not pointing at the sun: Direct CH1, NIP1, NIP2, NIP-RG780, LI201, TWC500, CUV4; and the following instruments were unshaded: 8-48, CM22, PSP, CUV4, PIR, CG4.
12BMSPYHR100Direct LI-20107/30/200912:45:0007/30/200923:59:00The Sci-Tec #2 tracker was apparently not working during this time. The following instruments were unshaded: CM22, CUV4, CG4; and the following instruments were not pointing at the sun: NIP #1, NIP RG780, LI-201.
13BMSDirect LI-20111/15/200814:02:0011/15/200814:32:00Due to a planned power outage at SRRL, the following BMS instruments were not pointing at the sun: Direct CH1, NIP1, NIP2, LI201, TWC500, CUV4.
14BMSPYHR100Direct LI-20105/25/200610:50:0005/25/200611:26:00Sun Follower III tracker was down due to power shutdown for installation of new SRRL equipment lift. The following instruments were NOT pointing at the SUN: NIP #1, NIP RG780, CUVB2, LI-201, TWC (500nm).
15BMSPYHR100Direct LI-20105/08/200608:30:0005/08/200615:30:00The LICOR sensor was signal was erratic and very large (up to 6000 W/m^2). Cause is unknown, however there may have been patching for BORCAL 2006-01 occuring on that day.
16BMSPYHR100Direct LI-20111/05/200506:37:0011/07/200508:10:00Sun Follower III tracker was down due to a power outage on 11/4/2005 at 23:14 MST. The following instruments were NOT pointing at the SUN: NIP #1, NIP RG780, CUVB2, LI-201.
17BMSPYHR100Direct LI-20107/23/200510:16:0007/24/200509:01:00Sun Follower III tracker was down due to a scheduled NREL power outage on 7/23/2005 at 5:33 MST. The following instruments were NOT pointing at the SUN: NIP #1, NIP RG780, CUVB2, LI-201.
18BMSPYHR100Direct LI-20107/16/200500:00:0007/18/200507:46:00The Sun Follower III tracker was hung up on cables, due to a research NIP removal that occured earlier in the week. This instrument was not pointing at the sun during this time.
19BMSPYHR100Direct LI-20101/01/200400:00:0009/23/200814:00:00The LI-201 has been having a moisture leak that has been gradually getting worse over time. Instrument was removed from service on 9/3/2008 @ 13:50 MST so that it could be disassembled and dried. Listed start date is approximate. Instrument was returned to service 9/23/2008 @ 14:00 MST, however instrument needs to be re-calibrated, since the annual calibrations on the unit were resulting in calibration factors that were artificially high (to compensate for moisture leak).